OpenStudy (peterg2001):

help please need help fast To solve -6x + 4 = -20, what steps would you perform? 4. -6.

OpenStudy (peterg2001):

@ThisIsAdriAtHerBest

OpenStudy (anonymous):

-6x + 4 = -20 -20 - 4 = ? ? / 6 = x

OpenStudy (anonymous):

Move the 4 to the other side by subtracting it from both sides, and then divide both sides by -6.

OpenStudy (anonymous):

*-6

OpenStudy (peterg2001):

anouther one please

OpenStudy (peterg2001):

To solve -9 - 16m = 7, what steps would you use? 9 -16

OpenStudy (anonymous):

Kind of the same question so, try it yourself.

OpenStudy (peterg2001):

answer please

OpenStudy (anonymous):

-9 - 16m = 7 7 + -9 = ? ? / 16 = m